# DySurv: Dynamic Deep Learning Model for Survival Analysis with Conditional Variational Inference

### v2

@@ II-C Metrics
In this section, we will switch the notation of samples from superscript to subscript, hence $\mathbf{x}^{i}$ is now $\mathbf{x}_{i}$ for sample $i$. Since we are no longer making single risk predictions at specific times and estimating distributions of event times for censored samples, different evaluation metrics must apply than those used in classic machine learning classification and prediction settings. The most common metric for evaluating survival analysis models is the concordance index $C_{ind}$, which estimates the probability that, for a random pair of samples, the predicted survival times (risk probabilities) of the two samples have the same ordering as their true survival times [5]. This explanation works perfectly for settings of proportional hazards where the ordering does not change over time but for our purposes, we will not be limited by such assumptions. Hence, we will rely on using the time-dependent extension $C_{ind}^{\mathrm{td}}$ with some modifications accounting for predictions independent of feature observations having a concordance of 0.5. The metric can be represented as
$C_{ind}^{\mathrm{td}}=\mathrm{P}\left\{\hat{S}\left(T_{i}\mid\mathbf{x}_{i}\right)<\hat{S}\left(T_{i}\mid\mathbf{x}_{j}\right)\mid T_{i}<T_{j},D_{i}=1\right\}$ (8)
@@
where $\hat{S}$ indicates the estimated survival probabilities are used and $D_{i}=1$ that only those who experienced the event are considered in this metric. A noted limitation of this metric is its obvious bias and dependence on the censoring distribution as only non-censored samples are considered making it affected by the length of stay and the censoring proportion that increases over the length of stay. To this end, we decided to use additional metrics for more holistic evaluation especially as previously proposed models like DeepHit were found to have lackluster results in real-world datasets when evaluated using other metrics besides concordance. We also evaluate our model using the Integrated Brier Score or IBS. The Brier Score is similar to the mean squared error as it represents the average squared distances between the predicted and the true survival probability (approximated with step functions with jumps at the event times) and is always a number between 0 and 1, with 0 being the best possible value [29]. The expectation of the Brier Score contains the mean squared error as one of its additive terms, so minimising one is minimising the other [20]. Since we need to know the event times for calculating IBS and we do not have access to all the samples’ event times in right-censoring, an adjusted metric called the inverse probability of censoring weights Brier Score (IPCW) is used instead to approximate the times by weighting the scores of the observed event times by the inverse probability of censoring. The equation used is thus
$\operatorname{BS}_{\operatorname{IPCW}}(t)=\frac{1}{n}\sum_{i=1}^{n}[\frac{\hat{S}_{i}(t)^{2}1\left\{T_{i}\leq t,D_{i}=1\right\}}{\hat{G}_{i}\left(T_{i}-\right)}+$ (9)
$\frac{\left[1-\hat{S}_{i}(t)\right]^{2}1\left\{T_{i}>t\right\}}{\hat{G}_{i}(t)}]$
@@
where $\hat{G}_{i}(t)=\mathrm{P}\left(C_{i}^{*}>t\right)>0$ is the Kaplan-Meier estimate of the censoring distribution for sample $i$ and $C_{i}^{*}$ is the censoring time. The expected value of this metric is the same as that for the uncensored Brier Score. As one notices, this metric is evaluated at specific times, whereas the Integrated Brier Score or IBS provides a general evaluation of model performance at all times.
@@
A limitation of this metric, however, is the biased assumption of the censoring distribution being the same across samples thereby disregarding covariate effects. This can be addressed by using an administrative extension of the metric that requires access to all the censoring times but a discussion of this in greater detail can be perused here [20]. Lastly, we introduce the IPCW (negative) binomial log-likelihood or NBLL from classic binary classification which measures both discrimination and calibration of the estimates and uses its integrated extension for all times INBLL
$\operatorname{BLL}(t)=\frac{1}{N}\sum_{i=1}^{N}[\frac{\log\left[1-\hat{S}\left(t\mid\mathbf{x}_{i}\right)\right]1\left\{T_{i}\leq t,D_{i}=1\right\}}{\hat{G}\left(T_{i}\right)}+$ (11)
$\frac{\log\left[\hat{S}\left(t\mid\mathbf{x}_{i}\right)\right]1\left\{T_{i}>t\right\}}{\hat{G}(t)}]$
@@ III Results
In survival analysis, it is also vital to show the performance of the models in creating survival curves, or estimates of the survival probabilities over time for different patients. Whereas traditional survival analysis models from statistics rely on risk sets and computing survival estimates for a population, an advantage of deep learning models is that the task of survival estimation can apply to an individual patient sample. We provide survival curves for a random group of five samples/patients (with different event times) across all datasets that show a clear separation of risk as can be seen in Figure 2. For MIMIC-IV, the data input consists of both static and time-series data and we provide survival curve results for both scenarios as well as for Dynamic-DeepHit for comparison. DySurv has a better capacity to separate independent patient risk trajectories as shown in Figure 3.
(a) Survival curve (estimate of survival probability over time) for five random samples from SUPPORT
(b) Survival curve (estimate of survival probability over time) for five random samples from METABRIC
(c) Survival curve (estimate of survival probability over time) for five random samples from GBSG
(d) Survival curve (estimate of survival probability over time) for five random samples from NWTCO
(e) Survival curve (estimate of survival probability over time) for five random samples from sac3
(f) Survival curve (estimate of survival probability over time) for five random samples from sac5
Figure 2: The set of survival curves for benchmark datasets as generated by DySurv shows clear extrapolation of risk across different samples. Since DySurv provides discrete estimates over time-intervals, the curves were generated with additional interpolation commonly used in other discrete methods in survival analysis.
@@
(c) Survival curve (estimate of survival probability over time) for five random samples from MIMIC-IV using both static and time-series features with Dynamic-DeepHit
Figure 3: The set of survival curves for the MIMIC-IV ICU EHR dataset as generated by DySurv that shows clear extrapolation of risk across different patients as compared to both static, time-series, and Dynamic-DeepHit cases. The same set of patients was used in generating the survival curves across cases.
@@
Figure 4: Comparison of DySurv at 24-hour prediction with existing ICU survival scores using (a) AUROC and (b) Sensitivity. APACHE IV was only able to be retrieved from MIMIC-IV and SOFA score was calculated from the eICU dataset. The calculations are available in the code repository.
IV Discussion
@@
As for MIMIC-IV and eICU results, we see that even with only static data in the input, DySurv manages to outperform other survival analysis models. When time-series data is added to the input in a multi-modal fashion, it can help the model improve its performance and outperform both CoxTime and Dynamic-DeepHit across all metrics. The survival curves in Figure 3 show clearly that the static version of DySurv has the clearest separation of survival trajectories for different patients and that the addition of time-series data makes the task more complex. This is probably due to the reconstruction task being a lot more difficult now that time-series are involved and the model makes sacrifices on the survival task front. However, compared to Dynamic-DeepHit, whose survival curves in Figure 3 c) are inseparable, the time-series version of DySurv still provides a relatively reliable survival estimate. We also see that, as we would expect clinically, the survival of the patients significantly changes in the last few days in the ICU, starting to drop a few days before death. Previous work has shown that earlier times in the ICU correspond to higher survival rates. This suggests that identifying the time period when survival rates drop dramatically in the ICU can help target earlier treatment for those most at risk [31].
V Conclusion
@@
In this paper, we present a novel dynamic risk prediction model for survival analysis based on deep learning in survival analysis paradigms. Our method builds on a combination of previous work including Dynamic-DeepHit by leveraging direct learning of the joint distribution of the first event time and the event through log-likelihood optimisation with logistic hazards. Theoretically, this approach is an alternative to the risk log-likelihood loss function of Dynamic-DeepHit itself that does not use a ranking loss for biased inflation of concordance results. Our DySurv model is capable of learning from complex EHR ICU time-series data and extracting lower-dimensional latent representations that can be useful for learning the survival task while also balancing reconstruction. As the model has been difficult to train due to loss instabilities and sensitivity to hyperparameter selection, future work can explore including a regularisation component to the loss terms. By using the underlying latent distribution, we can directly model an alternative to the survival distribution like the Weibull distribution instead of using a Gaussian intermediate.
VI Conflicts of interest
The authors declare no competing interests.
VII Code and Data sharing
The data used in the study are all publicly available. MIMIC-IV can be accessed after some initial requests and approvals from the PhysioNet website. The remaining datasets have their sources listed in the PyCox library repository. All of the code will be made available in a public GitHub repository after publication, including easily accessible notebooks to reproduce experiments from the paper.
VIII Acknowledgements
@@
Munib Mesinovic appreciates the support of the Rhodes Trust and the EPSRC Centre for Doctoral Training in Health Data Science. T.Z. was supported by the Royal Academy of Engineering under the Research Fellowship scheme.
Appendix A Appendix
A-A Data Description
A detailed list of features used in the study from different datasets can be seen in Tables V, VI, VII, VIII and IX.
TABLE V: Features extracted from the SUPPORT survival dataset.
Feature Type Feature Type
Sex binary Age integer
Number of comorbidities integer SUPPORT Coma Score continuous
Years of Education continuous APS III Score continuous
Diabetes binary Dementia binary
3rd Day MAP continuous 3rd Day WBC continuous
3rd Day HR continuous 3rd Day RR continuous
Temperature continuous 3rd Day Bilirubin continuous
TABLE VI: Features extracted from the NWTCO survival dataset.
Feature Type Feature Type
Relapse binary Age integer
Histology Local binary Histology Institute binary
Tumor Stage categorical Study Type categorical
TABLE VII: Features extracted from the GBSG survival dataset.
Feature Type Feature Type
Hormone Treatment binary Age integer
Tumor Grade categorical Menopause Treatment binary
Progesterone (fmol) continuous Number of positive nodes categorical
Estrogen (fmol) continuous
TABLE VIII: Features extracted from the METABRIC survival dataset.
Feature Type Feature Type
Age integer Gene I binary
Gene II binary Gene III binary
Hormone Treatment binary Gene IV binary
Radiotherapy binary Chemotherapy binary
ER-positive binary
TABLE IX: Features extracted from the MIMIC-IV database. The features include demographic data collected for all patients, ICU unit-specific information like the type of unit, hospital information like regional location, time since admission, vital signs including respiratory rate and blood pressure, and biochemical measurements including blood glucose and haemoglobin.
Static Variables
Feature Type Feature Type
Sex binary Motor Response continuous
Age integer Verbal Response continuous
Height continuous Ethnicity categorical
Weight continuous Unit Type categorical
Hour of Admission integer Admission Location categorical
Time Since Admission continuous
Time-series Variables
Feature Type Feature Type
GCS - Eye continuous
Eye Response continuous GCS - Motor continuous
Braden Score continuous GCS - Verbal continuous
Strength L Arm continuous Daily Weight continuous
Strength R Arm continuous ALT continuous
Strength L Leg continuous AST continuous
Strength R Leg continuous HCO3 continuous
Insurance categorical Hct continuous
ALT continuous Alkaline Phosphatase continuous
Anion Gap continuous AST continuous
Base Excess continuous Bicarbonate continuous
Bilirubin continuous Calcium continuous
Total CO2 continuous Chloride continuous
Creatinine continuous Glucose continuous
Hematocrit continuous Hemoglobin continuous
INR(PT) continuous Lactate continuous
MCH continuous MCHC continuous
MCV continuous Magnesium continuous
PT continuous PTT continuous
Phosphate continuous Platelet Count continuous
Potassium continuous RDW continuous
Red Blood Cells continuous Sodium continuous
Urea Nitrogen continuous White Blood Cells continuous
pCO2 continuous pH continuous
pO2 continuous JH-HLM continuous
Dyspnea Assessment continuous Daily Weight continuous
Glucose continuous Heart Rate continuous
DBP continuous SBP continuous
O2 Flow continuous O2 Sat (%) continuous
Pain Level continuous Pain Level Response continuous
Phosphorous continuous Respiratory Rate continuous
Richmond-RAS Scale continuous Temperature (${}^{\circ}$) continuous
TABLE X: Features extracted from the eICU database. The features include demographic data collected for all patients, ICU unit-specific information like type and number of beds, hospital information like regional location and teaching status, vital signs including respiratory rate and blood pressure, and biochemical measurements including troponin and levels of potassium and protein in the blood.
Feature Type Feature Type
Sex binary Unit Stay Type categorical
Age integer Num Beds Category categorical
Height continuous Region categorical
Weight continuous Teaching Status binary
Ethnicity categorical Physician Speciality categorical
Unit Type categorical Unit Type categorical
Unit Admit Source categorical Mechanical Ventilation binary
Unit Visit Number categorical
Time-series Variables
Feature Type Feature Type
Base Excess continuous
-basos continuous FiO2 continuous
-eos continuous HCO3 continuous
-monos continuous Hct continuous
-polys continuous Hgb continuous
ALT continuous MCH continuous
AST continuous MCHC continuous
BUN continuous MCV continuous
O2 Sat (%) continuous MPV continuous
PT-INR continuous PT continuous
RBC continuous PTT continuous
RDW continuous WBC continuous
Alkaline ph. continuous Albumin continuous
Bedside Glucose continuous Anion Gap continuous
Calcium continuous Bicarbonate continuous
Creatinine continuous Glucose continuous
Lactate continuous Magnesium continuous
pH continuous paCO2 continuous
paO2 continuous Phosphate continuous
Platelets continuous Potassium continuous
Sodium continuous Bilirubin continuous
Protein continuous Troponin - I continuous
Urinary s. Gravity continuous mean BP continuous
SBP continuous DBP continuous
A-B Benchmark Models
To compare DySurv to other survival analysis methods, we implement a selection of the most popular and consistently cited methods in the field and provide a short description of each below. We will first introduce discrete-time methods which rely on discretising the event times into specified durations, and then follow with continuous-time methods.
A-B1 PMF
The parametrisation of the Probability Mass Function (PMF) of the event times is another way of estimation without resorting to using discrete-time risk or hazard in likelihood optimisation. We described the continuous probability density function without using discrete time boundaries. It is the foundation of other methods like DeepHit and Multi-Task Logistic Regression. It similarly resorts to optimising a negative log-likelihood loss but instead of using the cumulative risk function, it uses the approximations of the PMF and survival functions [19]. Since we can establish a direct representational relation between the risk and survival functions, the PMF loss can be seen as an alternative to our loss function. In our and others’ implementation, the PMF method is a simple Multi-Layer Perceptron (MLP) optimised for this loss and we use the same structure as we use for our survival module to be described later.
A-B2 MTLR
Multi-Task Logistic Regression provides a generalization of the binomial log-likelihood to jointly model the sequence of binary labels for each time interval risk prediction. This method similarly minimises the negative log-likelihood with the PMF and survival function terms but the network outputs are cumulatively summed in reverse to no certain advantage and, in fact, just add computational complexity and numeric instability [27, 29].
A-B3 BCESurv
This Binary Cross-Entropy for Survival is a method consisting of a set of binary classifiers that remove individuals as they are censored. The loss is the binary cross entropy of the survival estimates at a set of discrete times, with targets that are indicators of surviving each time. Each output node in the last layer corresponds to a binary classifier evaluated at that time point. As censored patients are removed, the method is biased towards those with higher event probabilities [20].
A-B4 DeepHit
The single-risk version of DeepHit is a deep learning model whose output nodes are softmaxed to jointly model the probabilities between the event times and the time durations are discretised like in our case. The model depends on optimising both the negative log-likelihood loss based on the cumulative incidence function and a ranking loss built on the intuition of the concordance. The ranking loss penalises incorrect ordering of patient pairs in which the patient that remains longer in the study should have a lower risk at the endpoint for the patient with the shorter stay. Including this loss function allows the model to optimise for concordance. The concordance was the only evaluation metric used in the paper which might have lead to potentially biased and inflated results. Subsequently reproduced work has shown that indeed this model is not calibrated well and significantly lags across other metrics when compared to simpler models [7, 10, 23, 32].
A-B5 Logistic Hazard
The Logistic Hazard method is a submodular implementation of our deep learning model using the loss in (4) with an MLP that similarly parametrises the PMF of the survival times [28]. DySurv expands on this method to include it as a component in the framework with the variational autoencoder to jointly optimise for both tasks of reconstruction and latent space formation as well as survival estimation. A key difference between the log-likelihood loss used here (and in our model) and in DeepHit is that logistic hazards do not allow for survival past the maximum time horizon.
A-B6 CoxTime
CoxTime is a relative risk model that extends Cox regression beyond the proportional hazards and is the first of the continuous-time methods. The standard Cox regression model which we will not spend space introducing here consists of a baseline hazard term (defined cumulatively in the loss by a pre-selected estimator such as Breslow) and a relative risk term which is an exponential factor of the weighted linear combination of features. The basic model assumes constant proportionality between the patients’ hazards over time and is thus restrictive. In other words, the difference between survival likelihoods for a given time is proportional to the difference in feature or hazard values for patients. CoxTime goes around this assumption by parametrising the relative risk term as a function of time and not just the features, thus the non-proportional behaviour over time is modelled by allowing for time to be considered alongside the features [5].
A-B7 CoxCC
CoxCC (Case Control) is just a proportional implementation of the CoxTime model and is the closest to the standard Cox implementation where the minimisation of the partial log-likelihood is done with stochastic gradient descent by averaging over constrained risk sets for mini-batch learning instead of the entire dataset like in classical survival analysis [5].
A-B8 DeepSurv
DeepSurv is a deep learning model that directly minimises the negative partial log-likelihood as defined in the standard Cox model. It is one of the first deep learning implementations for survival analysis and the output of the model is the log-risk term of the Cox model which accounts for nonlinearity [6]. There is no indirect estimation of cumulative risk or survival through likelihood estimation like in the previous methods, thus DeepSurv similarly suffers under limitations of the Cox such as the proportionality assumption [33].
A-B9 PCHazard
The last continuous-time method we introduce and implement is PCHazard which assumes that the continuous-time hazard rate (instantaneous value of risk) is piecewise constant. The method relies on optimising for the likelihood contribution which mimics the MTLR approach albeit in the continuous-time setting with the hazards parametrised by a simple MLP. The piecewise constant causes the likelihood to behave like a Poisson likelihood [34]. Despite the method operating in continuous-time, the hazards are defined in time intervals which rely on discretisation steps from the observed continuous event times and censoring times, while we discretise the times to a predefined set of time flagposts [19].
A-C Model architecture
Following the LSTM unit, the remaining part of the encoder consists of an MLP module with 3 layers. The encoder and decoder are mirrored in their structure with the hidden neurons in the MLP layers consisting of 3 times feature length, then 5 times, then 3 times, before passing onto the last output layer. For the encoder, the last layer is used in approximating the mean and standard deviation of the latent Gaussian distribution. As for the decoder, the last output layer is used for reconstructing the input vector. The survival module is similar to the MLP components of the encoder and decoder. All components are jointly optimised through multi-loss optimisation. All of the components have been investigated with and without dropout included. The output of the survival module is 10 nodes softmaxed hence properly jointly distributed, each of which gives the probability that the patient has suffered the event (death) at that specific time interval.
Figure 5: Proposed DySurv model for dynamic risk prediction based on a combination of static and time-series EHR data
A-D Ablation studies
To empirically arrive at the optimal configuration for the architecture, we undertook ablation studies and grid search combinations for hyperparameters using a validation set presented in the tables below. The hyperparameters investigated for DySurv include:
  • Dropout: 0.5, 0.6, 0.7, 0.8
  • Number of hidden layers: 2, 3, 4, 5
  • Number of hidden neurons: 10, 25, 75, 100, current
  • Latent representation vector length: 10, 20, 30
  • Learning rate: 0.00001, 0.0001, 0.001, 0.01, 0.1
  • Batch size: 8, 16, 32, 64, 128, 256, 512
TABLE XI: Ablation study experiments with DySurv on MIMIC-IV using time-series and static components with the optimal hyperparameter combination. The top strategy was kept as the optimal one when adding lower strategies. For example, once the inclusion of dropout was found to contribute positively to the results, it was included in the model architecture, then the addition of batch normalisation was tested.
$C_{ind}^{\mathrm{td}}$ IBS IBLL
Dropout ($p$ = 0.5)
With 57.9 0.122 0.320
Without 53.1 0.145 0.389
Batch Normalisation
With 57.6 0.130 0.322
Without 57.9 0.122 0.320
Conditional term concatenation
With 57.9 0.122 0.320
Without 55.7 0.144 0.377
DySurv ($+$ time-series) 57.9 0.122 0.320

### v3

@@ II-C Metrics
In this section, we will switch the notation of samples from superscript to subscript, hence $\mathbf{x}^{i}$ is now $\mathbf{x}_{i}$ for sample $i$. Since we are no longer making single risk predictions at specific times and estimating distributions of event times for censored samples, different evaluation metrics must apply than those used in classic machine learning classification and prediction settings. The most common metric for evaluating survival analysis models is the concordance index $C_{ind}$, which estimates the probability that, for a random pair of samples, the predicted survival times (risk probabilities) of the two samples have the same ordering as their true survival times [5]. This explanation works perfectly for settings of proportional hazards where the ordering does not change over time but for our purposes, we will not be limited by such assumptions. Hence, we will rely on using the time-dependent extension $C_{ind}^{\mathrm{td}}$ with some modifications accounting for predictions independent of feature observations having a concordance of 0.5. The metric was originally proposed in [38] and all of the included metrics are provided in the PyCox package. The metric can be represented as
$C_{ind}^{\mathrm{td}}=\mathrm{P}\left\{\hat{S}\left(T^{i}\mid\mathbf{x}^{i}\right)<\hat{S}\left(T^{i}\mid\mathbf{x}^{j}\right)\mid T^{i}<T^{j},y^{i}=1\right\}$ (9)
@@
where $\hat{S}$ indicates the estimated survival probabilities are used and $y^{i}=1$ that only those who experienced the event are considered in this metric. A noted limitation of this metric is its obvious bias and dependence on the censoring distribution as only non-censored samples are considered making it affected by the length of stay and the censoring proportion that increases over the length of stay. To this end, we decided to use additional metrics for more holistic evaluation especially as previously proposed models like DeepHit were found to have lackluster results in real-world datasets when evaluated using other metrics besides concordance. We also evaluate our model using the Integrated Brier Score or IBS. The Brier Score is similar to the mean squared error as it represents the average squared distances between the predicted and the true survival probability (approximated with step functions with jumps at the event times) and is always a number between 0 and 1, with 0 being the best possible value [39]. The expectation of the Brier Score contains the mean squared error as one of its additive terms, so minimising one is minimising the other [22]. Since we need to know the event times for calculating IBS and we do not have access to all the samples’ event times in right-censoring, an adjusted metric called the inverse probability of censoring weights Brier Score (IPCW) is used instead to approximate the times by weighting the scores of the observed event times by the inverse probability of censoring. The equation used is thus
$\operatorname{BS}_{\operatorname{IPCW}}(t)=\frac{1}{n}\sum_{i=1}^{n}[\frac{\hat{S}^{i}(t)^{2}\mathbb{1}\left\{T^{i}\leq t,y^{i}=1\right\}}{\hat{G}^{i}\left(T^{i}-\right)}+$ (10)
$\frac{\left[1-\hat{S}^{i}(t)\right]^{2}\mathbb{1}\left\{T^{i}>t\right\}}{\hat{G}^{i}(t)}]$
@@
where $\hat{G}^{i}(t)=\mathrm{P}\left(C^{i*}>t\right)>0$ is the Kaplan-Meier estimate of the censoring distribution for sample $i$ and $C^{i*}$ is the censoring time. The expected value of this metric is the same as that for the uncensored Brier Score. As one notices, this metric is evaluated at specific times, whereas the Integrated Brier Score or IBS provides a general evaluation of model performance at all times.
@@
A limitation of this metric, however, is the biased assumption of the censoring distribution being the same across samples thereby disregarding covariate effects. This can be addressed by using an administrative extension of the metric that requires access to all the censoring times but a discussion of this in greater detail can be perused where the metric was originally proposed [22]. Lastly, we introduce the IPCW (negative) binomial log-likelihood or NBLL from classic binary classification from [22] which measures both discrimination and calibration of the estimates and uses its integrated extension, INBLL, for all times
$\operatorname{BLL}(t)=\frac{1}{N}\sum_{i=1}^{N}[\frac{\log\left[1-\hat{S}\left(t\mid\mathbf{x}^{i}\right)\right]1\left\{T^{i}\leq t,y^{i}=1\right\}}{\hat{G}\left(T^{i}\right)}+$ (12)
$\frac{\log\left[\hat{S}\left(t\mid\mathbf{x}^{i}\right)\right]1\left\{T^{i}>t\right\}}{\hat{G}(t)}]$
@@ III RESULTS
In survival analysis, it is also vital to show the performance of the models in creating survival curves, or estimates of the survival probabilities over time for different patients. Whereas traditional survival analysis models from statistics rely on risk sets and computing survival estimates for a population, an advantage of deep learning models is that the task of survival estimation can apply to an individual patient sample. We provide survival curves for a random group of five samples/patients (with different event times) across all datasets that show a clear separation of risk as can be seen in Figure 2. For MIMIC-IV, the data input consists of both static and time-series data and we provide survival curve results for both scenarios in Figure 3.
(a) Survival curve for five random samples from SUPPORT
(b) Survival curve for five random samples from METABRIC
(c) Survival curve for five random samples from GBSG
(d) Survival curve for five random samples from NWTCO
(e) Survival curve for five random samples from sac3
(f) Survival curve for five random samples from sac5
Figure 2: Survival curves (estimate of survival probability over time) for benchmark datasets by DySurv across different samples. DySurv provides discrete estimates over time and additional interpolation was applied. Dots correspond to true event times which were predicted correctly by DySurv.
@@
Figure 3: The set of survival curves for the MIMIC-IV ICU EHR dataset as generated by DySurv shows extrapolation of risk across different patients as compared to static and time-series feature sets. Dots correspond to true event times which were predicted correctly by DySurv.
@@
Figure 4: Comparison of DySurv at 24-hour prediction with existing ICU survival scores and deep learning survival models on MIMIC-IV using (a) AUROC and (b) Sensitivity and c) AUPRC. APACHE IV was only able to be retrieved from MIMIC-IV and SOFA score was calculated from the eICU dataset.
IV CONCLUSIONS
@@
As for MIMIC-IV and eICU results, we see that even with only static data in the input, DySurv manages to outperform other survival analysis models. When time-series data are added to the input in a multi-modal fashion, it can help the model improve its performance and outperform both CoxTime and Dynamic-DeepHit across all metrics. Figure 3 shows examples of projected survival trajectories using DySurv in both the static-only and static-and-time-series feature settings. We also see that, as expected in the hospitalised setting, the survival of the patients significantly changes in the last few days in the ICU, starting to drop a few days before death. Previous work has shown that earlier discharge times in the ICU correspond to higher survival rates. This suggests that identifying the period when survival rates drop dramatically in the ICU can help target earlier treatment for those most at risk [41]. We assume that patients will stay in the ICU for a maximum of 10 days and while that is generally true, there could be outliers and differences in ICU stay duration distributions. We suspect this would not severely hinder the generalizability of DySurv to those scenarios as that would just mean an expansion of the time component for the survival trajectory predictions.
From Figure 4 we can see how DySurv outperforms existing in-use systems for patient risk prognosis in the ICU across a variety of metrics. We acknowledge the limitation of the results for generalisability from using different data subsets. The model is highly dependent on the quality and completeness of time-series data. While forward and backward filling has been commonly employed and verified with clinical collaborators [26, 25], this approach could introduce biases, particularly if the missingness is non-random or systematically tied to patient outcomes. Furthermore, we used pre-defined time intervals (72-hour windows) which may not fully capture the dynamic variability in patient trajectories, particularly for patients with more erratic health progressions. This time interval, however, can be manually adjusted depending on the identified application. We also acknowledge potential overfitting to the benchmark datasets, especially given the relatively small size of some of these datasets. This could hinder generalization to more diverse populations or clinical settings where patient characteristics and ICU protocols differ significantly. We have found that careful hyperparameter optimisation can aid in addressing this challenge when learning on different datasets.
The generalizability of DySurv is also constrained by the assumptions embedded in the model architecture, despite being nonparametric. While DySurv does not assume proportional hazards, it still relies on the assumptions inherent in its logistic hazards model producing discrete estimates of the survival probabilities. Additionally, as deep learning models are inherently opaque, their extensions to healthcare settings come with concerns of trustworthiness and interpretability. To address this limitation and further verify that DySurv is learning relevant clinical patterns from the patient data, we used permutation importance to obtain the top 10 predictive features for DySurv on MIMIC-IV and eICU. These results are in line with relevant clinical literature and practice and a discussion can be found in Section G in the Supplementary.
@@
In this paper, we present a novel dynamic risk prediction model for survival analysis based on deep learning in survival analysis paradigms. We work collaboratively with intensive care physicians and demonstrate interdisciplinary pathways to the implementation of deep learning for survival analysis methods in the ICU as shown in Figure 2 in the Supplementary. Our method builds on a combination of previous work including Dynamic-DeepHit by leveraging direct learning of the joint distribution of the first event time and the event through log-likelihood optimisation with logistic hazards. Theoretically, this approach is an alternative to the risk log-likelihood loss function of Dynamic-DeepHit itself, which does not use a ranking loss for biased inflation of concordance results. Our DySurv model is capable of learning from complex EHR ICU time-series data and extracting lower-dimensional latent representations that can be useful for learning the survival task while also balancing reconstruction. As the model has been difficult to train due to loss instabilities and sensitivity to hyperparameter selection, future work can explore including a regularization component to the loss terms. By using the underlying latent distribution, we can directly model an alternative to the survival distribution like the Weibull distribution instead of using a Gaussian intermediate.
V Funding Statement
@@
This work was supported by the Rhodes Trust and the EPSRC Centre for Doctoral Training in Health Data Science grant (EP/S02428X/1).
VI Competing Interests Statement
The authors have no competing interests to declare.
VII Contributorship Statement
MM, PW, and TZ have all contributed to the conception or design of the work; or the acquisition, analysis, or interpretation of data for the work. MM has extracted and pre-processed the data with guidance from PW and TZ. MM has developed the method with guidance from TZ, designed the experiments, validated the results, and summarised the findings. MM, PW, and TZ drafted the work or reviewed it critically for important intellectual content. MM provided the initial draft and implemented the subsequent changes in writing. MM, PW, and TZ approved the final version for publication. All authors agree to be accountable for all aspects of the work in ensuring that questions related to the accuracy or integrity of any part of the work are appropriately investigated and resolved.
VIII DATA AVAILABILITY
The data is available from public access requests for eICU and MIMIC-IV. Accessing the data requires ethics module training and certification. It can be obtained at the following links: https://physionet.org/content/eicu-crd/2.0/ and https://physionet.org/content/mimiciv/2.0/.